slipknot3066:

The equation of a line is −6x−2y=−18. What is the x-intercept of the line?

AngeI:

First put the equation in y=mx+b -6x - 2y = -18 -2y = 6x + 18 y = -3x - 9 Now, substitute 0 for x to find the x-intercept y = -3(0) - 9 y = -9 X-intercept = (0, -9)

slipknot3066:

omg thanks